When acceleration be a function of velocity as a= f(v).. Then :-
a) The displacement (x) = integral vdv/f(v)
b) the acceleration may be constant
c) the slope of acceleration versus velocity graph may may be constant.
d) a) & c) are correct

Please give solution

Grade:12

1 Answers

Nirmal Singh.
10 years ago
a = f(x)

dv / dt = f(x)

(dx / dt)*(dv/ dx) = f(x)
v dv / dx = f(x)
so dx = vdv / f(x)

hence x = integral of (vdv) / f (x)

Hence Answer a is correct and

acceleration is directly proportionlal to velocity so acce. will not be constant

a = f(x)

so a / f(x) = 1

hencethe slope of acceleration versus velocity graph may may be constant

Hence Answer a and c is correct

Regards,
Nirmal Singh
Askiitians Faculty